# Faithful-First Reasoning, Planning, and Acting for Multimodal LLMs

### v3

@@ Abstract
Multimodal Large Language Models (MLLMs) frequently suffer from unfaithfulness, generating reasoning chains that drift from visual evidence or contradict final predictions. We propose Faithful-First Reasoning, Planning, and Acting (RPA) framework in which FaithEvi provides step-wise and chain-level supervision by evaluating the faithfulness of intermediate reasoning, and FaithAct uses these signals to plan and execute faithfulness-aware actions during inference. Experiments across multiple multimodal reasoning benchmarks show that faithful-first RPA improves perceptual faithfulness by up to 24% over prompt-based and tool-augmented reasoning frameworks, without degrading task accuracy. Our analysis shows that treating faithfulness as a guiding principle perceptually faithful reasoning trajectories and mitigates hallucination behavior. This work thereby establishes a unified framework for both evaluating and enforcing faithfulness in multimodal reasoning. Code will be released upon acceptance.
@@
Junxian Li1thanks: Equal contribution.   Xinyue Xu211footnotemark: 1   Sai Ma311footnotemark: 1   Di Zhang4   Sichao Li5thanks: Correspondence: scli@cityu.edu.mo 1Shanghai Jiao Tong University 2The Hong Kong University of Science and Technology 3The Australian National University 4 Fudan University 5City University of Macau
@@ 1 Introduction
Despite rapid progress in multimodal large language models (MLLMs) Sun et al. (2023); Chen et al. (2024b); An et al. (2025b); Bai et al. (2025), their reasoning traces remain unfaithful: models frequently produce persuasive explanations that conflict with perceptual evidence, or utilize post-hoc rationalizations fabricate their reasoning progress Arcuschin et al. (2025); Barez et al. (2025). This gap poses a central challenge for trustworthy reasoning. Existing efforts typically focus on improving task accuracy or enriching Chain-of-Thought (CoT) generation Zhang et al. (2023), yet the unfaithful reasoning remains unaddressed.
@@ 2 Perceptual and Behavioral Faithfulness
Previous studies mainly concern whether a reasoning trace accurately reflects the decision process that led to the model’s final output Lyu et al. (2023); Matton et al. (2025); Barez et al. (2025); Arcuschin et al. (2025), which we refer to as Behavioral faithfulness (BF). While important, it underestimates the importance of input evidence in the reasoning process. We therefore propose Perceptual faithfulness (PF), which concerns whether individual reasoning steps are grounded in verifiable input evidence, such as entities and attributes that are perceptually present in the input. A perceptually unfaithful explanation may invoke objects and properties that are unsupported or contradicted by the visual input. Importantly, the model is not required to output the correct answer to be faithful (Jacovi and Goldberg, 2020; Dasgupta et al., 2022).
@@ 3.1 Extraction of Claimed Objects
Concretely, we query Qwen2.5-7B-Instruct (Qwen)  Team (2024) with a structured prompt (see Appx. B) to extract, for each step $s_{t}$, a set $O_{t}$ of $m_{t}$ claimed objects, where $O_{t}=\{O_{t}^{1},O_{t}^{2},\dots,O_{t}^{m_{t}}\}.$ The union of all extracted objects across the reasoning chain is denoted as
@@ 3.2.2 Grounding
For each object $O_{t}^{i}$ we employ a frozen GroundingDINO-base model Liu et al. (2023b) to detect and ground the object in the input image $I$. GroundingDINO returns a set of bounding boxes and associated confidence scores:
@@ 3.3 Faithfulness Scoring
$f_{t}^{i}=\begin{cases}0,&c_{t}^{i}<0.4,\quad\text{(object confidently absent)}\\ c_{t}^{i},&0.4\leq c_{t}^{i}\leq 0.6,\quad\text{(uncertain existence)}\\ 1,&c_{t}^{i}>0.6,\quad\text{(object confidently present)}.\end{cases}$
@@ 5.1 Experimental Setup
Datasets. We evaluate our framework on widely used multimodal benchmarks covering object recognition, visual grounding, and hallucination-sensitive question answering, including LLaVA-bench Liu et al. (2023a), RealWorldQA xAI (2024), POPE Li et al. (2023b), and MMHal-Bench Sun et al. (2023) (MMHal). These datasets feature images with rich real-world objects and are well suited for evaluating perceptual faithfulness. Specifically, the queried objects in POPE are removed to ensure fair assessment.
Baselines. We compare our framework with representative training-free reasoning frameworks, including CoT Wei et al. (2022); Zhang et al. (2023), ReAct Yao et al. (2022), Grounded-CoT Wu et al. (2025a), and Visual Abstract Thinking (VAT) Liu et al. (2025). These methods are evaluated on widely used MLLMs, namely Qwen-2.5-VL-7B (Qwen) Bai et al. (2025), InternVL3-8B (InternVL) Zhu et al. (2025), and LLaVA-one-Vision-1.5-8B (LLaVA) An et al. (2025b).
@@ Appendix A Related Work
Faithfulness in Multimodal Reasoning. Despite the demonstrated effectiveness of reasoning in improving the performance of MLLMs, an unfaithful problem emerges: the reasoning traces generated by these models do not always behave faithfully to the internal processes that produce their final answers Lyu et al. (2023); Turpin et al. (2023); Lanham et al. (2023). Models often rely on latent knowledge or shortcut associations that are not explicitly expressed in their reasoning chains. As a result, the generated reasoning steps may read as a plausible but untrustworthy explanation, motivating the need for explicit faithfulness assessment Jacovi and Goldberg (2020); Parcalabescu and Frank (2023); Matton et al. (2025); Barez et al. (2025); Arcuschin et al. (2025). Several works have introduced benchmarks or metrics to evaluate faithfulness explicitly. M4 Li et al. (2023a), FaithScore Jing et al. (2023) and TIFA Hu et al. (2023) propose metrics for evaluating the faithfulness for vision-language models. However, we find that behavioral alignment (behavioral faithfully) does not guarantee the correctness of the final output (see left panel in Fig.1).

### v4

@@ Abstract
Multimodal Large Language Models (MLLMs) frequently suffer from unfaithfulness, generating reasoning chains that drift from visual evidence or contradict final predictions. We propose Faithful-First Reasoning, Planning, and Acting (RPA) framework in which FaithEvi provides step-wise and chain-level supervision by evaluating the faithfulness of intermediate reasoning, and FaithAct uses these signals to plan and execute faithfulness-aware actions during inference. Experiments across multiple multimodal reasoning benchmarks show that faithful-first RPA improves perceptual faithfulness by up to 24% over prompt-based and tool-augmented reasoning frameworks, without degrading task accuracy. Our analysis shows that treating faithfulness as a guiding principle perceptually faithful reasoning trajectories and mitigates hallucination behavior. This work thereby establishes a unified framework for both evaluating and enforcing faithfulness in multimodal reasoning. Code is at https://github.com/lijunxian111/Faithful-First-RPA.
@@
Junxian Li1thanks: Equal contribution.   Xinyue Xu211footnotemark: 1   Sai Ma311footnotemark: 1   Di Zhang4   Sichao Li5thanks: Correspondence: sichao.li@sydney.edu.au 1Shanghai Jiao Tong University 2The Hong Kong University of Science and Technology 3The Australian National University 4 Fudan University 5University of Sydney lijunxian0531@sjtu.edu.cn, xinyue.xu@connect.ust.hk, sai.ma@anu.edu.au di.zhang@ustc.edu, sichao.li@sydney.edu.au
@@ 1 Introduction
Despite rapid progress in multimodal large language models (MLLMs) Sun et al. (2023); Chen et al. (2024b); An et al. (2025b); Wang et al. (2024), their reasoning traces remain unfaithful: models frequently produce persuasive explanations that conflict with perceptual evidence, or utilize post-hoc rationalizations fabricate their reasoning progress Arcuschin et al. (2025); Barez et al. (2025). This gap poses a central challenge for trustworthy reasoning. Existing efforts typically focus on improving task accuracy or enriching Chain-of-Thought (CoT) generation Zhang et al. (2023), yet the unfaithful reasoning remains unaddressed.
@@ 2 Perceptual and Behavioral Faithfulness
Previous studies mainly concern whether a reasoning trace accurately reflects the decision process that led to the model’s final output Lyu et al. (2023); Matton et al. (2025); Barez et al. (2025); Arcuschin et al. (2025); Li et al. (2026); Ma et al. (2026), which we refer to as Behavioral faithfulness (BF). While important, it underestimates the importance of input evidence in the reasoning process. We therefore propose Perceptual faithfulness (PF), which concerns whether individual reasoning steps are grounded in verifiable input evidence, such as entities and attributes that are perceptually present in the input. A perceptually unfaithful explanation may invoke objects and properties that are unsupported or contradicted by the visual input. Importantly, the model is not required to output the correct answer to be faithful (Jacovi and Goldberg, 2020; Dasgupta et al., 2022).
@@ 3.1 Extraction of Claimed Objects
Concretely, we query Qwen2.5-7B-Instruct (Qwen)  Yang et al. (2024) with a structured prompt (see Appx. B) to extract, for each step $s_{t}$, a set $O_{t}$ of $m_{t}$ claimed objects, where $O_{t}=\{O_{t}^{1},O_{t}^{2},\dots,O_{t}^{m_{t}}\}.$ The union of all extracted objects across the reasoning chain is denoted as
@@ 3.2.2 Grounding
For each object $O_{t}^{i}$ we employ a frozen GroundingDINO-base model Liu et al. (2024) to detect and ground the object in the input image $I$. GroundingDINO returns a set of bounding boxes and associated confidence scores:
@@ 3.3 Faithfulness Scoring
$f_{t}^{i}=\begin{cases}0,&c_{t}^{i}<0.4,\quad\text{(confidently absent)}\\ c_{t}^{i},&0.4\leq c_{t}^{i}\leq 0.6,\text{(uncertain existence)}\\ 1,&c_{t}^{i}>0.6,\quad\text{(confidently present)}.\end{cases}$
@@ 5.1 Experimental Setup
Datasets. We evaluate our framework on widely used multimodal benchmarks covering object recognition, visual grounding, and hallucination-sensitive question answering, including LLaVA-bench Liu et al. (2023), RealWorldQA xAI (2024), POPE Li et al. (2023b), and MMHal-Bench Sun et al. (2023) (MMHal). These datasets feature images with rich real-world objects and are well suited for evaluating perceptual faithfulness. Specifically, the queried objects in POPE are removed to ensure fair assessment.
Baselines. We compare our framework with representative training-free reasoning frameworks, including CoT Wei et al. (2022); Zhang et al. (2023), ReAct Yao et al. (2022), Grounded-CoT Wu et al. (2025a), and Visual Abstract Thinking (VAT) Liu et al. (2025). These methods are evaluated on widely used MLLMs, namely Qwen-2.5-VL-7B (Qwen) Wang et al. (2024), InternVL3-8B (InternVL) Zhu et al. (2025), and LLaVA-one-Vision-1.5-8B (LLaVA) An et al. (2025b).
@@ 5.4 Ablation and Sensitivity Analysis
we also choose Qwen on RealWorldQA for an additional ablation analysis over various threshold values. Results are reported in Table 4. Notably, the changes in hyperparameters do not influence the final performance of FaithAct too much. The threshold 0.5-0.5 will lead to a slight decline in performance.
@@
$\alpha$ & threshold 0.3-0.7 0.4-0.6 0.5-0.5
[…]
Table 4: Sensitivity analysis on $\alpha$ and threshold.
5.5 Computational Overhead
We report the average time per chain generation (in seconds) using Qwen-2.5-VL-7B under identical decoding settings in Table 5. Importantly, our method does not modify the underlying generation function or decoding procedure of the MLLM. Consequently, the token generation rate (tokens/second) remains effectively the same as that of raw CoT. We also report reasoning-chain lengths in Appendix H.
5.6 Perception-light Benchmarks
we have additionally evaluated FaithAct on MathVista Lu et al. (2024) and ScienceQA Lu et al. (2022), where visual evidence is necessary but not sufficient and abstract or external knowledge is often required. Table 6 reports the $F_{\text{chain}}$ of FaithAct and raw CoT (%) on these datasets. FaithAct performs well on ScienceQA and slightly better than CoT on MathVista.
6 Discussion: Behavioral Faithfulness
To provide more analysis and insights between behavioral faithfulness and perceptual faithfulness, we added a qualitative taxonomy to showing that perceptual faithfulness (PF) and behavioral faithfulness (BF) can dissociate across all four regimes (BF+ PF+; BF+ PF-; BF- PF+; BF- PF-, ’+’ and ’-’ mean faithful and unfaithful). Importantly, PF+ does not guarantee BF+, where a model may correctly ground intermediate observations but still produce an incorrect final decision due to downstream reasoning errors. Conversely, BF+ does not imply PF+, where a model can arrive at the correct answer while relying on perception-unfaithful rationales. We will also clarify a standard operationalization of behavioral faithfulness from prior works in the main text. For example, if the model truly uses the cited evidence, then removing or perturbing that evidence should systematically alter the model’s decision.
@@ 7 Conclusion
Datasets LLaVA-BENCH REALWORLDQA MMHAL POPE
FaithAct (s) 18.85 14.04 16.48 14.72
CoT (s) 10.73 4.65 5.59 3.41
$\delta$ Time 8.12 9.39 10.89 11.31
Table 5: Inference time comparison.
Dataset & Methods MathVista ScienceQA
[…]
Table 6: Faithfulness evaluation on perception-light benchmarks.
@@
Acknowledgements
This work was supported in part by computing resources provided by Lambda AI.
@@ Appendix A Related Work
Faithfulness in Multimodal Reasoning. Despite the demonstrated effectiveness of reasoning in improving the performance of MLLMs, an unfaithful problem emerges: the reasoning traces generated by these models do not always behave faithfully to the internal processes that produce their final answers Lyu et al. (2023); Turpin et al. (2023); Lanham et al. (2023). Models often rely on latent knowledge or shortcut associations that are not explicitly expressed in their reasoning chains. As a result, the generated reasoning steps may read as a plausible but untrustworthy explanation, motivating the need for explicit faithfulness assessment Jacovi and Goldberg (2020); Parcalabescu and Frank (2023); Matton et al. (2025); Barez et al. (2025); Arcuschin et al. (2025); Peng et al. (2025a, b). Several works have introduced benchmarks or metrics to evaluate faithfulness explicitly. M4 Li et al. (2023a), FaithScore Jing et al. (2023) and TIFA Hu et al. (2023) propose metrics for evaluating the faithfulness for vision-language models. However, we find that behavioral alignment (behavioral faithfully) does not guarantee the correctness of the final output (see left panel in Fig.1).
@@ G.1 Study Design and Data Collection
Dataset & Metrics LLaVA-bench RealWorldQA MMHal POPE
Avg token length (FaithAct) 270.88 171.43 112.85 149.11
Avg token length (CoT) 357.52 162.78 200.06 148.23
Min token length (FaithAct) 30 17 12 28
Min token length (CoT) 98 64 45 43
Max token length (FaithAct) 512 512 510 512
Max token length (CoT) 512 432 502 512
Table 7: Token length statistics for FaithAct and CoT across different benchmarks.
@@ G.2 Evaluation Metrics
Appendix H Reasoning-Chain Length
We report the reasoning chain length of our method of Qwen-2.5-VL-7B in Table 7, compared with raw CoT. The official tokenizer of the model is used. Here we report the metrics calculated from the final output of FaithAct. We find that FaithAct does not produce longer reasoning chains than raw CoT. In fact, on certain metrics, it yields shorter chains. A plausible explanation is that FaithAct constrains the model to reason only over visually verified objects with acceptable faithfulness scores. By filtering out unsupported or hallucinated objects, the framework thereby leads to more concise and focused chains.